# Robust Bernoulli mixture models for credit portfolio risk

### v1

@@ 4.3 Robustness Results for siBMMs
$\mathcal{F}_{icx}^{\alpha}:=\{f\in\mathcal{F}_{icx}\mid f(0)=0\,,f(1)=\alpha\,,||f||_{L}\leq 1\}$ (4.5)
of increasing convex functions mapping from $(0,0)$ to $(1,\alpha)$ with Lipschitz constant $||\cdot||_{L}$ being not larger than $1\,.$ It is clear that $\mathcal{F}_{icx}^{\alpha}$ is closed under minima and maxima, i.e., for an arbitrary family $(f_{t})_{t\in I}\subset\mathcal{F}_{icx}^{\alpha}\,,$ also the pointwise minimum and maximum
$\bigwedge_{t\in I}f_{t}(x)$ $:=\min_{t\in[0,1]}\{f_{t}(x)\}\,,$ $\bigvee_{t\in I}f_{t}(x)$ $:=\max_{t\in[0,1]}\{f_{t}(x)\}$ (4.6)
are increasing convex functions satisfying the same boundary and Lipschitz conditions. We make use of the function class $\mathcal{F}_{icx}^{\alpha}$ to describes default integral function as defined in (2.2). It can easily be seen that any default integral function associated with an siBBM is in the class $\mathcal{F}_{icx}^{\alpha}$ for $\alpha=\pi_{n}\,.$ Vice versa, for every function $f\in\mathcal{F}_{icx}^{\alpha}\,,$ there exists a siBMM with $\pi_{n}=\alpha$ having default integral function $G_{D_{n},Z}=f\,.$
The following theorem addresses classes of siBMMs and determines improved lower and upper bounds in convex order for the associated credit portfolio losses. These improvements are relative to the standard bounds for sums of positively dependent random variables within the same Fréchet class. In the context of credit portfolio models, the standard bounds are given by the independent loss $L^{\perp}$ and the comonotonic loss $L^{c}$ as defined in (4.3).
@@
For fixed default probabilities $\pi=(\pi_{1},\ldots,\pi_{N})\,,$ let $\{(D_{1}^{x},\ldots,D_{N}^{x},Z^{x})\}_{x\in I}$ be a family of siBMMs with $(D_{1}^{x},\ldots,D_{N}^{x},Z^{x})\in\mathcal{M}^{\pi}$, with default integral functions $G_{D^{x}_{n},Z^{x}}$, and with losses $L^{x}:=\sum_{n=1}^{N}e_{n}\delta_{n}D_{n}^{x}$ for $x\in I\,.$ Consider siBMMs models $(\underline{D}_{1},\ldots,\underline{D}_{N},\underline{Z})$ and $(\overline{D}_{1},\ldots,\overline{D}_{N},\overline{Z})$ in $\mathcal{M}^{\pi}$ specified by the default integral functions $\underline{G}_{n}$ and $\overline{G}_{n}$ defined by
$\underline{G}_{n}(s)$ $:=\bigvee_{x\in I}G_{D_{n}^{x},Z^{x}}(s)\quad\text{and}\quad\overline{G}_{n}(s):=\bigwedge_{x\in I}G_{D_{n}^{x},Z^{x}}(s)\,$ (4.7)
for all $s\in[0,1]\,.$ Then, we have
@@
    Theorem 4.5 imposes no assumption on the mixing variables $Z^{x}$. Instead, the bounds are derived solely based on the pointwise minima and maxima of the default integral functions, ensuring broad applicability and minimal structural constraints. The functions $\underline{G}_{n}$ and $\overline{G}_{n}$ again define siBMMs, which yield the bounds $\underline{L}$ and $\overline{L}$ for $L^{x}$ in convex order.
@@
    The lower bound $\underline{L}$ (upper bound $\overline{L}$) in Theorem 4.5 is sharp if there exists a sequence $(x_{m})_{m\in\mathbb{N}}\,,$ $x_{m}\in I$ for all $m\,,$ of siBMMs such that $G_{D_{n}^{x_{m}},Z^{x_{m}}}(s)\to\underline{G}_{n}(s)$ ($G_{D_{n}^{x_{m}},Z^{x_{m}}}(s)\to\overline{G}_{n}(s)$) for all $s$ as $m\to\infty\,.$ In this case, the derivatives of the convex default integral functions converge pointwise (outside a Lebesgue-null set) so that the assertion follows from the dominated convergence theorem. The conditions are satisfied, for instance, in a parametric subfamily of Gaussian or Clayton copula-based threshold models, where the underlying parameters lie within specified intervals; see also Examples 3.6 and 4.2.
@@
    If in a BMM $(D_{1},\ldots,D_{N},Z)$ the risk factor $Z$ is independent from the default variables $D_{1},\ldots,D_{N}\,,$ then we have $P(D_{n}=1\mid Z=F_{Z}^{-1}(t))=\pi_{n}$ and thus $G_{n,Z}(t)=t\pi_{n}$ for all $t\in[0,1]$ and for all $n\,.$ Hence, the lower bound $\underline{L}$ improves the loss $L^{\perp}$ of the independence model in convex order if and only if there exists $n\in\{1,\ldots,N\}$ and $s\in(0,1)$ such that $\underline{G}_{n,Z}(s)\neq s\pi_{n}\,,$ see also Remark 2.2 (e). Similarly, if $D_{1},\ldots,D_{N},Z$ are comonotonic, then $P(D_{n}=1\mid Z=F_{Z}^{-1}(t))=\mathds{1}_{\{t\geq 1-\pi_{n}\}}$ and thus $G_{n,Z}(t)=(t-1+\pi_{n})\mathds{1}_{\{t\geq 1-\pi_{n}\}}$ for all $t$ and for all $n\,.$ Hence, the upper bound $\overline{L}$ improves the comonotonic bound $L^{c}$ in convex order if and only if there exists $n\in\{1\ldots,N\}$ and $t\in(0,1)$ such that $\overline{G}_{n,Z}(t)\neq(t-1+\pi_{n})\mathds{1}_{\{t\geq 1-\pi_{n}\}}\,.$
@@
    For a family =$\mathcal{F}=\{f_{t}\}_{t\in I}$ of functions in $\mathcal{F}_{icx}^{\alpha}\,,$ denote by
    $\mathfrak{c}(\mathcal{F}):=\left\{f\in\mathcal{F}_{icx}^{\alpha}\mid\bigwedge_{t\in I}f_{t}(s)\leq f(s)\leq\bigvee_{t\in I}f_{t}(s)\leavevmode\nobreak\ \text{for all }s\in(0,1)\right\}$ (4.9)
    the class of increasing convex functions generated by $\mathcal{F}\,.$ It can easily be seen that the functions $g_{1}(s)=\alpha s$ and $g_{2}(s)=(s-1+\alpha)\mathds{1}_{\{s\geq 1-\alpha\}}$ generate the class $\mathcal{F}_{icx}^{\alpha}\,,$ i.e., $\mathfrak{c}(\{g_{1},g_{2}\})=\mathcal{F}_{icx}^{\alpha}\,.$ The right-hand plot in Figure 2.1 illustrates for $\alpha=0.2$ the classes $\mathfrak{c}(\{g_{1},g_{2}\})$ and $\mathfrak{c}(\{h_{1},h_{2}\})$ as the area between the respective functions, where $h_{1}(s)=1/12s^{2}+7/60s$ and $h_{2}(s)=0.2s^{3}\,.$
To quantify portfolio losses in terms of risk measures, we consider convex, law-invariant risk measures defined on the space of integrable random variables on $(\Omega,\mathcal{A},P)\,.$ It is well known that convex law-invariant risk measures are consistent with the convex order if the underlying probability space is atomless, see Bäuerle and Müller (2006); Burgert and Rüschendorf (2006); Jouini et al. (2006). For an overview of risk measures, we refer to (Föllmer and Schied, 2016, Chapter 4). The following result is a direct consequence of Theorem 4.5.
@@
The AVaR is a benchmark risk measure in the Basel III and Solvency III framework for market risk. It was introduced as a replacement for VaR in order to better capture tail risks in the revised approach to risk measurement implemented in Basel Committee on Banking Supervision (2013). Note that the VaR is not a convex risk measure, but is still used as the benchmark under the Internal Ratings Based (IRB) approach of the Basel regulatory framework for the calculation of minimum capital requirements for credit portfolios. In the following section, we apply our robustness results to data and study improved bounds for the AVaR for families of siBMMs under parameter uncertainty and for classes of siBMMs under model uncertainty.
@@ 5.1 Simulation study
First, we perform a simulation study considering four different classes of threshold models. The first one, the Gaussian copula model, is a standard copula model but has the drawback that it does not admit tail dependencies. Threshold models, as discussed in Section 3, define the default indicator by the event where the value of a financial asset drops below a specified threshold, typically corresponding to the company’s liability value. Given the strong tail dependencies commonly observed in financial assets, copula models that account for such dependencies are generally more appropriate than the Gaussian copula model. For this reason, we specifically examine the Clayton copula family, which belongs to the class of Archimedean copulas, along with its survival copulas. As a fourth class, we examine a hybrid approach that combines Gaussian and Clayton copula models. We then compare the bounds derived under model uncertainty for this union with the individual bounds obtained under parameter uncertainty for each copula model.
@@
Our analysis is based on an application of Corollary 4.7 in order to determine uncertainty intervals for the AVaR of credit portfolio losses in the respective classes of siBBMs. To this end, we need to compare the associated default integral functions $G_{D_{n},Z}$ pointwise. If the distribution function of the common risk factor $Z$ is continuous [^4] If $F_{Z}$ is discontinuous, then the conditional default probability $P(D_{n}=1\mid Z=F_{Z}^{-1}(s))$ can be similarly expressed as a generalized partial derivative of the underlying copula $C\,,$ see (Ansari and Rüschendorf, 2021, Theorem 2.2). Then the representation of $G_{D_{n},Z}$ in (5.1) has to be modified slightly., $G_{D_{n},Z}$ has a simple representation through the copula of $(D_{n},Z)$ as
$\begin{split}G_{D_{n},Z}(t)&=\int_{0}^{t}P(D_{n}=1\mid Z=F_{Z}^{-1}(s))\mathrm{\,d}s\\ &=\int_{0}^{t}[1-P(D_{n}=0\mid Z=F_{Z}^{-1}(s))]\mathrm{\,d}s\\ &=t-\int_{0}^{t}\partial_{2}C_{D_{n},Z}(1-\pi_{n},F_{Z}(F_{Z}^{-1}(s)))\mathrm{\,d}s\\ &=t-C_{D_{n},Z}(1-\pi_{n},t)\,,\end{split}$ (5.1)
where the first equality is given by the definition of $G_{n,Z}$ in (2.2). The third equality follows from the representation of the conditional distribution function in (3.6) and the transformation $V=F_{Z}(Z)$ using continuity of $F_{Z}\,.$ The last equality follows from $F_{Z}(F_{Z}^{-1}(s))=s$ for all $s$ using again continuity of $F_{Z}\,.$ Hence, due to (5.1), a pointwise comparison of the default integral functions $G_{D_{n},Z}$ follows from a pointwise comparison of the underlying copulas in the Bernoulli mixture models.
Since, by Proposition 3.3 and Lemma 3.5, every siBMM $(D_{1},\ldots,D_{N},Z)$ can be identified with a copula-based threshold model $D_{n}=\mathds{1}_{\{X_{n}\leq c_{n}\}}$, where the copula of $C_{X_{n},Y}$ is SI, $n\in\{1,\ldots,N\}$, we obtain from (3.8) the representation
$G_{D_{n},Z}(s)=s-\hat{C}_{X_{n},Y}(1-\pi_{n},s)$ (5.2)
@@
In the following example, we consider copula-based threshold models, where the dependence structure of $(X_{n},Y)$ is described by a Gaussian copula, a Clayton copula, and a survival Clayton copula, respectively. In each case, the underlying copula parameter is subject to uncertainty.
Example 5.1 (Bernoulli mixture models under parameter uncertainty).
We consider three siBMMs implied by copula-based threshold models under parameter uncertainty and apply Corollary 4.7 to obtain the minimal and maximal credit portfolio risks for the respective family of models. The probability of default is set to $\pi_{n}=0.02$ for all $n\in\{1,\ldots,N\}\,.$
@@
    Gaussian copula model: For all $n\in\{1,\ldots,N\}\,,$ we consider a family $\{D_{n}^{\rho_{n}}=\mathds{1}_{\{X_{n}\leq c_{n}\}}\}_{\rho_{n}\in I_{n}}$ of threshold models, where $X_{n}$ follows the Merton model (3.4) with unspecified asset correlation $\rho_{n}\in I_{n}:=[0.12,0.24]$ for all $n\in\{1,\ldots,N\}\,.$ The bounds of the interval $I_{n}$ for $\rho_{n}$ are chosen as the standard bounds $0.12$ and $0.24$ in the IRB approach (see Basel Committee on Banking Supervision (2005) or Basel Committee on Banking Supervision (2011)), where the asset correlation is determined by
@@
    i.e. as a function of the borrower’s unconditional default probability $\pi_{n}$. The joint distribution of $X_{n}$ and the latent risk factor $Y$ is bivariate normal with correlation $\sqrt{\rho_{n}}\,.$ Hence, the associated copula is the Gaussian copula with parameter $\sqrt{\rho_{n}}\,.$ Due to (3.3), the threshold is given as $c_{n}=\Phi^{-1}(\pi_{n})\,.$ According to Example 3.6 (a), the implied family of Bernoulli mixture models $\{(D_{1}^{\rho_{1}},\ldots,D_{N}^{\rho_{N}},Z)\}_{\rho_{1},\ldots,\rho_{N}}$ is specified through the copulas $C_{D_{n}^{\rho_{n}},Z}=C_{\sqrt{\rho_{n}}}^{Ga}\,.$ Due to (5.2) the associated default integral functions are given by
    $\begin{split}G_{D_{n}^{\rho_{n}},Z}(s)=s-\widehat{C}_{X_{n},Y}(1-\pi_{n},s)&=s-\widehat{C_{\sqrt{\rho_{n}}}^{Ga}}(1-\pi_{n},s)\\ &=s-\Phi\left(\frac{\Phi^{-1}(1-\pi_{n})-\sqrt{\rho_{n}}\,\Phi^{-1}(s)}{\sqrt{1-\rho_{n}}}\right)\,.\end{split}$ (5.4)
    Since the Gaussian copula family $(C^{Ga}_{\rho})_{\rho\in[-1,1]}$ is increasing in $\rho$ with respect to the pointwise order of copulas and since Gaussian copulas with non-negative parameter are SI, we obtain that the default integral bounds specified in Theorem 4.5 are given by
    $\underline{G}^{Ga}_{n}(s)$ $:=\bigvee_{\rho_{n}\in I}G_{D_{n}^{\rho_{n}},Z}(s)=G_{D_{n}^{0.24},Z}(s)=s-\Phi\left(\frac{\Phi^{-1}(1-\pi_{n})-\sqrt{0.24}\,\Phi^{-1}(s)}{\sqrt{1-0.24}}\right)\,,$ (5.5)
    $\overline{G}^{Ga}_{n}(s)$ $:=\bigwedge_{\rho_{n}\in I}G_{D_{n}^{\rho_{n}},Z}(s)=G_{D_{n}^{0.12},Z}(s)=s-\Phi\left(\frac{\Phi^{-1}(1-\pi_{n})-\sqrt{0.12}\,\Phi^{-1}(s)}{\sqrt{1-0.12}}\right)\,,$ (5.6)
    see also Example 4.6(b). The upper left plot in Figure 5.1 illustrates the range of the default integral functions in the Gaussian threshold model under parameter uncertainty, where, for all $n\in\{1,\ldots,N\}\,,$ the upper (lower) dashed graph corresponds to the lower (upper) default integral bound $\underline{G}^{Ga}_{n}$ ($\overline{G}^{Ga}_{n}$). The solid line indicates the default integral function in the IRB approach for the asset correlation $\rho_{n}\approx 0.165$ obtained from formula (5.3) for our choice $\pi_{n}=0.02\,.$
@@
    Clayton copula model: Analogous to (a), we consider for each $n$ a family $\{D_{n}^{\theta_{n}}=\mathds{1}_{\{X_{n}\leq c_{n}\}}\}_{\theta_{n}\in I}$ of threshold models, where now $X_{n}$ depends on the latent factor $Y$ via the Clayton copula model (3.14) with unspecified dependence parameter $\theta_{n}\in I_{n}:=[\underline{\theta},\overline{\theta}].$ For reasons of comparability, the interval bounds $\underline{\theta}:=4\pi\arcsin(0.12)/(1-2\pi\arcsin(0.12))$ and $\overline{\theta}:=4\pi\arcsin(0.24)/(1-2\pi\arcsin(0.24))$ are chosen as the parameters of the Clayton copula having the same Kendall’s tau value as the Gaussian copula with the parameter bounds from the IRB formula as discussed in part (a). Since copulas have uniform marginals, the threshold level is given by $c_{n}=\pi_{n}\,.$ The default integral functions of the implied family $\{(D_{1}^{\theta_{1}},\ldots,D_{N}^{\theta_{N}},Z)\}_{\theta_{1},\ldots,\theta_{N}}$ of siBMMs are given by

### v2

@@ 3 Robustness Results for Credit Risk Models
[…]
of increasing convex functions mapping from $(0,0)$ to $(1,\alpha)$ with Lipschitz constant $||\cdot||_{L}$ being not larger than $1.$ For an arbitrary family $f=(f_{t})_{t\in I}\subset\mathcal{F}_{icx}^{\alpha},$ we denote by
$\overline{\mathop{co}}(f)$ $:=\inf\{g\mid g\penalty 10000\ \text{convex},g(y)\geq\sup_{t\in I}f_{t}(y)\penalty 10000\ \forall y\},$
$\underline{\mathop{co}}(f)$ $:=\sup\{g\mid g\penalty 10000\ \text{convex},g(y)\leq\inf_{t\in I}f_{t}(y)\penalty 10000\ \forall y\},$
the least convex majorant and greatest convex minorant. It is straightforward to verify that also $\underline{\mathop{co}}(f)$ and $\overline{\mathop{co}}(f)$ are in $\mathcal{F}_{icx}^{\alpha}.$ We make use of the function class $\mathcal{F}_{icx}^{\alpha}$ to describe default integral functions as defined in (2.2). It can easily be seen that any default integral function associated with an siBMM is in the class $\mathcal{F}_{icx}^{\alpha}$ for $\alpha=\pi_{n}.$ Vice versa, for every function $g\in\mathcal{F}_{icx}^{\alpha},$ there exists a siBMM with $\pi_{n}=\alpha$ having default integral function $G_{D_{n},Z}=g.$
The following theorem addresses classes of siBMMs and determines improved lower and upper bounds in convex order for the associated credit portfolio losses. These improvements are relative to the standard bounds for sums of positively dependent random variables within the same Fréchet class, given by the independent loss $L^{\perp}$ and the comonotonic loss $L^{c}$.
@@
For fixed default probabilities $\pi=(\pi_{1},\ldots,\pi_{N}),$ let $\{(D_{1}^{x},\ldots,D_{N}^{x},Z^{x})\}_{x\in I}$ be a family of siBMMs with $(D_{1}^{x},\ldots,D_{N}^{x},Z^{x})\in\mathcal{M}^{\pi}$, with default integral functions $G_{D^{x}_{n},Z^{x}}$, and with losses $L^{x}:=\sum_{n=1}^{N}e_{n}\delta_{n}D_{n}^{x}$ for all $x\in I.$ Consider siBMMs $(\underline{D}_{1},\ldots,\underline{D}_{N},\underline{Z})$ and $(\overline{D}_{1},\ldots,\overline{D}_{N},\overline{Z})$ in $\mathcal{M}^{\pi}$ specified by the default integral functions
$\underline{G}_{n}$ $:=\overline{\mathop{co}}\left((G_{D_{n}^{x},Z^{x}})_{x\in I}\right)\quad\text{and}\quad\overline{G}_{n}:=\underline{\mathop{co}}\left((G_{D_{n}^{x},Z^{x}})_{x\in I}\right),$ (3.3)
respectively. Then, we have
@@
    Theorem 3.1 imposes no assumption on the mixing variables $Z^{x}$. Instead, the bounds are derived solely based on the least convex majorant and greatest convex minorant of the family of default integral functions, ensuring broad applicability and minimal structural constraints. The functions $\underline{G}_{n}$ and $\overline{G}_{n}$ in (3.3) again define siBMMs, which yield the bounds $\underline{L}$ and $\overline{L}$ for $L^{x}$ in convex order. Note that $\underline{G}_{n}\geq\overline{G}_{n}$ pointwise, in accordance with Remark 2.2(a) where pointwise smaller default integral functions indicate stronger positive dependencies.
@@
    The lower bound $\underline{L}$ (upper bound $\overline{L}$) in Theorem 3.1 is sharp if there exists a sequence $(x_{m})_{m\in\mathbb{N}},$ $x_{m}\in I$ for all $m,$ of siBMMs such that $G_{D_{n}^{x_{m}},Z^{x_{m}}}(s)\to\underline{G}_{n}(s)$ ($G_{D_{n}^{x_{m}},Z^{x_{m}}}(s)\to\overline{G}_{n}(s)$) for all $s$ as $m\to\infty.$ In this case, the derivatives of the convex default integral functions converge pointwise (outside a Lebesgue-null set) so that the assertion follows from the dominated convergence theorem. The conditions are satisfied, for instance, in a parametric subfamily of Gaussian or Clayton copula-based threshold models, where the underlying parameters lie within specified intervals; see also Examples A.5 and A.7.
@@
    If in a BMM $(D_{1},\ldots,D_{N},Z)$ the risk factor $Z$ is independent of the default variables $D_{1},\ldots,D_{N},$ then we have $\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(t))=\pi_{n}$ and thus $G_{n,Z}(t)=t\pi_{n}$ for all $t\in[0,1]$ and for all $n.$ Hence, the lower bound $\underline{L}$ improves the loss $L^{\perp}$ of the independence model in convex order if and only if there exists $n\in\{1,\ldots,N\}$ and $s\in(0,1)$ such that $\underline{G}_{n,Z}(s)\neq s\pi_{n}.$ Similarly, if $D_{1},\ldots,D_{N},Z$ are comonotonic and $F_{Z}$ is continuous, then $\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(t))=\mathds{1}_{\{t\geq 1-\pi_{n}\}}$ and thus $G_{n,Z}(t)=(t-1+\pi_{n})\mathds{1}_{\{t\geq 1-\pi_{n}\}}$ for all $t$ and for all $n.$ Hence, the upper bound $\overline{L}$ improves the comonotonic bound $L^{c}$ in convex order if and only if there exists $n\in\{1\ldots,N\}$ and $t\in(0,1)$ such that $\overline{G}_{n,Z}(t)\neq(t-1+\pi_{n})\mathds{1}_{\{t\geq 1-\pi_{n}\}}.$
@@
    For a family $f=\{f_{t}\}_{t\in I}$ of functions in $\mathcal{F}_{icx}^{\alpha},$ denote by
    $\mathfrak{c}(f):=\left\{h\in\mathcal{F}_{icx}^{\alpha}\mid\underline{\mathop{co}}(f)(s)\leq h(s)\leq\overline{\mathop{co}}(f)(s)\penalty 10000\ \text{for all }s\in(0,1)\right\}$ (3.5)
    the class of increasing convex functions generated by $f.$ The functions $g_{1}(s)=\alpha s$ and $g_{2}(s)=(s-1+\alpha)\mathds{1}_{\{s\geq 1-\alpha\}}$ generate the class $\mathcal{F}_{icx}^{\alpha},$ i.e., $\mathfrak{c}(\{g_{1},g_{2}\})=\mathcal{F}_{icx}^{\alpha}.$ The right-hand plot in Figure 2.1 illustrates for $\alpha=0.2$ the classes $\mathfrak{c}(\{g_{1},g_{2}\})$ and $\mathfrak{c}(\{h_{1},h_{2}\})$ as the area between the respective functions, where $h_{1}(s)=1/12s^{2}+7/60s$ and $h_{2}(s)=0.2s^{3}.$
To quantify portfolio losses in terms of risk measures, we consider convex, law-invariant risk measures defined on the space of integrable random variables on $(\Omega,\mathcal{A},P).$ It is well known that convex law-invariant risk measures are consistent with the convex order if the underlying probability space is atomless; see Bäuerle and Müller (2006); Burgert and Rüschendorf (2006); Jouini et al. (2006). For an overview of risk measures, we refer to (Föllmer and Schied, 2016, Chapter 4). The following result is a direct consequence of Theorem 3.1.
@@
In the Basel III and Solvency II frameworks, the AVaR is a benchmark risk measure in the internal models for market risk. It was introduced as a replacement for VaR in order to better capture tail risks in the revised approach to risk measurement implemented in Basel Committee on Banking Supervision (2013). Note that the VaR is not a convex risk measure, but is still used as the benchmark under the Internal Ratings Based (IRB) approach of the Basel regulatory framework for the calculation of minimum capital requirements for credit portfolios. In the following section, we apply our robustness results to data and study improved bounds for the AVaR for families of siBMMs under parameter uncertainty and for classes of siBMMs under model uncertainty.
@@ 4.1 Simulation study
We perform a simulation study considering four different classes of threshold models. In threshold models, default is defined as the event where the value of a financial asset drops below a specified threshold, typically corresponding to the company’s liability value; see Appendix A. As a standard threshold model, we consider the Gaussian copula model which, however, has the drawback that it does not admit tail dependencies. Given the strong tail dependencies commonly observed in financial assets, copula models that account for such dependencies are generally preferable. For this reason, we specifically examine the Clayton copula family, which belongs to the class of Archimedean copulas, along with its survival copulas. As a fourth class, we examine a hybrid approach that combines Gaussian and Clayton copula models. We then compare the bounds derived under model uncertainty for this hybrid model with the individual bounds obtained for each copula model under parameter uncertainty.
@@
Our analysis is based on an application of Corollary 3.3 in order to determine uncertainty intervals for the AVaR of credit portfolio losses in the respective classes of siBBMs. To this end, we compare the associated default integral functions $G_{D_{n},Z}$ pointwise. By Proposition A.2 and Lemma A.4, every siBMM $(D_{1},\ldots,D_{N},Z)$ can be identified with a copula-based threshold model $D_{n}=\mathds{1}_{\{X_{n}\leq c_{n}\}}$, where the copula of $C_{X_{n},Y}$ is SI, $n\in\{1,\ldots,N\}$; see Appendix A.1 for all details. If the distribution function of the common risk factor $Z$ is continuous, we obtain from (A.7) the representation
$G_{D_{n},Z}(t)=\pi_{n}-C_{X_{n},Y}(\pi_{n},1-t)$ (4.1)
@@
In the following example, we consider copula-based threshold models, where the dependence structure of $(X_{n},Y)$ is described by a Gaussian, a Clayton, and a survival Clayton copula, respectively. In each case, the underlying copula parameter is subject to uncertainty.
Example 4.1 (siBMMs under parameter uncertainty).
We consider three siBMMs implied by copula-based threshold models (due to Proposition A.2) under parameter uncertainty and apply Corollary 3.3 to obtain the minimal and maximal credit portfolio risks for the respective family of models. The probability of default is set to $\pi_{n}=0.02$ for all $n\in\{1,\ldots,N\}.$
Figure 4.1: Default integral functions $G_{D_{n},Z}$ of the threshold model classes considered in Examples 4.1 and 4.2 under parameter and model uncertainty: Gaussian copula model (green), Clayton copula model (blue), and survival Clayton copula model (red). The solid lines are obtained from asset correlations estimated by the IRB formula, where the dashed lines $\overline{G}_{n}^{Ga}$, $\underline{G}_{n}^{Ga}$, $\overline{G}_{n}^{Cl}$, $\underline{G}_{n}^{Cl}$, $\overline{G}_{n}^{sCl}$, $\underline{G}_{n}^{sCl}$ correspond to the asset correlation bounds of the IRB formula (4.2).
@@
    Gaussian copula model: For $n\in\{1,\ldots,N\},$ we consider a family $\{D_{n}^{\rho_{n}}=\mathds{1}_{\{X_{n}\leq c_{n}\}}\}_{\rho_{n}\in I_{n}}$ of threshold models, where $X_{n}$ follows the Merton model (A.3) with unspecified asset correlation $\rho_{n}\in I_{n}:=[0.12,0.24]$. The bounds of the interval $I_{n}$ for $\rho_{n}$ are chosen as the standard bounds $0.12$ and $0.24$ in the IRB approach (see Basel Committee on Banking Supervision (2005, 2011)), where the asset correlation is determined as
@@
    which is a function of the borrower’s unconditional default probability $\pi_{n}$. The joint distribution of $X_{n}$ and the latent risk factor $Y$ is bivariate normal with correlation $\sqrt{\rho_{n}}.$ Hence, the associated copula is the Gaussian copula with parameter $\sqrt{\rho_{n}}.$ The threshold is given as $c_{n}=\Phi^{-1}(\pi_{n});$ see (A.2). According to (4.2) and Example A.5 (a), the implied family of siBMMs $\{(D_{1}^{\rho_{1}},\ldots,D_{N}^{\rho_{N}},Z)\}_{\rho_{1},\ldots,\rho_{N}}$ is specified by the default integral functions
    $\begin{split}G_{D_{n}^{\rho_{n}},Z}(s)=\pi_{n}-C_{\sqrt{\rho_{n}}}^{Ga}(\pi_{n},1-s)=\pi_{n}-\Phi_{\sqrt{\rho_{n}}}(\Phi^{-1}(\pi_{n}),\Phi^{-1}(1-s)),\end{split}$ (4.3)
    where $\Phi_{\rho}$ is the standardized bivariate normal cdf with correlation $\rho$ and $\Phi$ is the univariate standard normal cdf. Since the Gaussian copula family $(C^{Ga}_{\rho})_{\rho\in[-1,1]}$ is increasing in $\rho$ with respect to the pointwise order of copulas and since Gaussian copulas with non-negative parameter are SI, $G_{D_{n}^{\rho_{n}},Z}(s)$ is decreasing in $\rho_{n}$ in the pointwise order. Hence, the default integral bounds due to Theorem 3.1 are given by
    $\overline{G}^{Ga}_{n}(s)$ $:=\underline{\mathop{co}}\left((G_{D_{n}^{\rho_{n}},Z})_{\rho_{n}\in I})\right)(s)=G_{D_{n}^{0.24},Z}(s)=\pi_{n}-\Phi_{\sqrt{0.24}}(\Phi^{-1}(\pi_{n}),\Phi^{-1}(1-s)),$
    $\underline{G}^{Ga}_{n}(s)$ $:=\overline{\mathop{co}}\left((G_{D_{n}^{\rho_{n}},Z})_{\rho_{n}\in I})\right)(s)=G_{D_{n}^{0.12},Z}(s)=\pi_{n}-\Phi_{\sqrt{0.12}}(\Phi^{-1}(\pi_{n}),\Phi^{-1}(1-s)),$
    see also Remark 3.2(b). The upper left plot in Figure 4.1 illustrates the range of the default integral functions in the Gaussian threshold model under parameter uncertainty, where, for all $n\in\{1,\ldots,N\},$ the lower (upper) dashed graph corresponds to the default integral bound $\overline{G}^{Ga}_{n}$ ($\underline{G}^{Ga}_{n}$). The solid line indicates the default integral function in the IRB approach for the asset correlation $\rho_{n}\approx 0.165$ obtained from formula (4.2) for our choice $\pi_{n}=0.02.$
@@
    Clayton copula model: Analogous to (a), we consider for each $n$ a family $\{D_{n}^{\theta_{n}}=\mathds{1}_{\{X_{n}\leq c_{n}\}}\}_{\theta_{n}\in I}$ of threshold models, where now $X_{n}$ depends on the latent factor $Y$ via the Clayton copula model (A.9) with unspecified dependence parameter $\theta_{n}\in I_{n}:=[\underline{\theta},\overline{\theta}].$ For reasons of comparability, the interval bounds $\underline{\theta}:=4\pi\arcsin(0.12)/(1-2\pi\arcsin(0.12))$ and $\overline{\theta}:=4\pi\arcsin(0.24)/(1-2\pi\arcsin(0.24))$ are chosen as the parameters of the Clayton copula having the same Kendall’s tau value as the Gaussian copula with the parameter bounds from the IRB formula as discussed in part (a). Since copulas have uniform marginals, the threshold level is given by $c_{n}=\pi_{n}.$ The default integral functions of the implied family $\{(D_{1}^{\theta_{1}},\ldots,D_{N}^{\theta_{N}},Z)\}_{\theta_{1},\ldots,\theta_{N}}$ of siBMMs are given by